OpenStudy (anonymous):

equation: 2x + 3y = 1,200 Describe how you would graph this line using the slope-intercept method.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmm

OpenStudy (triciaal):

slope-intercept formula y = mx + b rearrange eqn in that format

OpenStudy (triciaal):

when x = 0 what is y? when y = 0 what is x?

OpenStudy (anonymous):

I don't get why you're replacing the x and y with 0?

OpenStudy (triciaal):

the easiest values to use to get 2 of the points to graph when x = 0 y = 1200 so you have the point (0, 1200) when y = 0 find x this is another point on the line connect the points to get the line satisfying y= -2/3x + 1200

OpenStudy (triciaal):

you can use any x values to calculate the corresponding y value and then get points on the line to graph

OpenStudy (jdoe0001):

recall the 400 was the y-intercept, thus is where the line touches the y-axis, or INTERCEPTS it thus the point (0, 400)

OpenStudy (jdoe0001):

if the slope is positive the graph goes UP from left to right if it's negative the graph goes DOWN from left to right lemme recorrect my correction darn =)

OpenStudy (anonymous):

so y=(0,400) and x is ??????????? sorry im just confused on what is plotted on the x axis.

OpenStudy (jdoe0001):

hold... on I got a few things mixed up

OpenStudy (triciaal):

correction the intercept is 1200/3 = 400 the point is (0, 400)

OpenStudy (triciaal):

@MonikaCristello do you have any questions about the method?

OpenStudy (jdoe0001):